OpenStudy (anonymous):

If the principal is $200 and the interest rate is 4 percent, what is the simple interest earned in one year? A.) $8 B.) $50 C.) $800

OpenStudy (anonymous):

M=C+Cdt M=200+(200)(0.04)(1) M=200+8 M=208$ M-C= Interest earned 208-200=8$

OpenStudy (anonymous):

If you need an explanation for it, shout me out, the variables are the variables I use in french as I study in french.

OpenStudy (anonymous):

i understand it the way that you put it (: thank you!

OpenStudy (anonymous):

No problem :)